OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

I could really use your help :P I suck at math, especially these type of problems.

OpenStudy (lucaz):

put the original line in slope intercept form y=3x+3 the slope is 3 (from 3x), so the slope of the parallel line is -1/3(negative reciprocal) as it passes through(4,-2) you can use the point slope form to find the equation of this line: \[\large{ y-y_1=m(x-x_1)\\ y-(-2)=-\frac{1}{3}(x-(4))\\ y+2=-\frac{x}{3}+\frac{4}{3}\\ y=-\frac{x}{3}-\frac{2}{3} }\]

OpenStudy (anonymous):

I think that I understand it...? Thank You :) @lucaz

OpenStudy (lucaz):

yw